# Examining Long-Context Large Language Models for Environmental Review Document Comprehension

### v2

@@ 3.1 Context Variation
No Context: In this baseline scenario, questions are presented to the models without any additional context. While this strategy can work well with popular domains such as sport or literature, we assume that the NEPA domain may be challenging for the LLM s to provide accurate answers. In other words, this setting can be considered a test of the LLM s’ ability to answer out-of-general-domain questions. As such, this setting is usually expected to return low performance.
@@
PDF Context: In this scenario, in addition to the question, we also provide the model with text contents from the entire PDF document, from which the benchmark question was generated. Since we do not inform the model which part of the document to look at, the accuracy of generated responses will heavily depend on the models’ ability to pick the correct context from the very large scale textual information provided. We expect this setting to yield performance better than the ‘no context’ scenario. This setup evaluates the models’ ability to process and extract relevant information from lengthy, complex documents.
@@
Augmented Context:  Employing a RAG approach, this setup provides the top-$k$ most relevant text chunks retrieved from a vector database containing text chunks of NEPA documents. In our RAG model implementation, when a question is inputted for LLM generation, the corresponding context is extracted as a relevant passage from a given NEPA document. We use the standard RAG setup where we encode both question and retrieved passages with the BGE embedding model Xiao et al. (2023). Cosine similarity scores are used to assess the similarity between the question and the contexts. In our case, the number of top-ranked relevant text chunks extracted is set at $k=3$. This scenario assesses the models’ performance when given highly relevant, focused information that is specifically retrieved to answer the question at hand.
@@
Gold Context:  In this configuration, we include the actual context from which the question was generated in the prompt, alongside the question content. This scenario represents an ideal case where the most relevant information is perfectly identified and provided. hile the situation where users manually identify the correct text passage is rare in practice, we simulate this scenario to measure how well LLM s can perform, if we were able to extract relevant passages with very high accuracy. This setup serves as an upper bound for performance, testing the models’ ability to comprehend and utilize perfectly relevant information.
3.2 Evaluation Metrics
To assess the performance of the models across different configurations, we employ a sophisticated evaluation approach that goes beyond simple syntactic similarity measures. While overlap based metrics such as BLEU scores Papineni et al. (2002) have been used in prior works, they are not ideal for evaluations where semantic accuracy is crucial. Therefore, we have chosen to use the RAGAS evaluation framework Es et al. (2023), specifically focusing on the answer correctness score. The answer correctness score combines two key aspects – (i) factual correctness: this measures the accuracy of the generated answer at the phrase/clause level compared to the ground truth. We utilize GPT-4 to quantify the factual overlap between the generated answer and the expected answer Es et al. (2023). (ii) semantic correctness: this is calculated by comparing the vector embeddings of the expected answers and the predicted answers using the BGE embedding model Xiao et al. (2023). In our evaluation, we assign different weights to these aspects – $75\%$ for factual correctness (weight of 0.75), and $25\%$ for semantic correctness (weight of 0.25). This weighting allows us to prioritize factual accuracy while still considering the overall semantic similarity of the responses. By using this comprehensive evaluation metric, we aim to provide a more nuanced and accurate assessment of the LLM s’ performance in handling NEPA-related queries across the different context scenarios. This approach enables us to capture both the precise factual content and the broader semantic accuracy of the generated responses, offering valuable insights into the models’ capabilities in processing and utilizing relevant document information.
@@ 3.2 Evaluation Metrics
We employ a separate evaluation method for closed type questions which have only ‘yes’ or ‘no’ answers using the following steps. First, we process the responses generated by the LLM s to extract only the ‘yes’ or ‘no’ answer from the generated response. Any additional explanatory text or elaboration is removed from the response. Thereafter, the cleaned binary response is compared directly with the ground truth answer. If the cleaned response exactly matches the ground truth, it is assigned a full score of $100\%$. If there is any discrepancy, including cases where the model fails to provide a clear "Yes" or "No" answer, the response is scored as $0\%$. This straightforward evaluation method for closed questions allows us to assess the models’ ability to provide accurate binary responses without being influenced by any uncertainties introduced by the RAGAS evaluation framework. We refer this metric as answer correctness score for closed type questions.
4 Performance Analysis
In this section, we describe the overall performance of the LLM s in the question/answering task evaluated with NEPAQuAD1.0 (as presented in Section 2). First, we compare the performance of five long-context LLM s: GPT-4, Claude-3 Sonnet, Gemini, Mistral and Llama3.1 across various QA contexts (Section 4.1). Second, we compare the model performance across various question types (Section 4.2). Finally, we analyze the performance.
4.1 Evaluating Different QA Contexts
@@
Table 2 reports the overall performance of the models across various QA contexts used in the evaluation. We observed that for the task with no context, the GPT-4 model produces the most accurate results by far. However, when PDF documents are provided as context, Claude surpassing GPT-4 and Gemini in the answer correctness. Despite that Gemini is able to handle very long contexts (1.5M tokens), it is surprising to see its performance drop when provided with PDF documents as additional contexts. This may be due to the model struggling to reason over the large amount of relevant and irrelevant content in the EIS document.
Open-source models (LLama3.1 and Mistral) generally underperform compared to closed LLM s across most context types, including PDF, RAG, and gold contexts. However, in the no-context scenario, all models exhibit similar performance, suggesting that the availability and quality of context significantly influence the performance gap between open-source and closed models.
@@
Table 2: Answer correctness of LLM responses on the NEPAQuAD benchmark across different context types. The best-performing setting for each model is shown in bold.
@@
4.2 Evaluating Different Question Types
@@
Figure 3: The evaluation results meassured by the Answer Correctness scores of each LLM used with 4 scenarios of using context over each question types
@@
In this section, we summarize few interesting observations when analyzing the performance of the models across different question types. The ‘process’, ‘comparison’ and ‘evaluation’ questions generally benefit from richer contexts, with closed-source models demonstrating a more significant improvement. ‘Convergent’ and ‘recall’ type questions prove challenging for all models, especially in no-context scenarios, highlighting the difficulty of synthesizing information or retrieving specific details without supporting context. ‘Problem-solving’ questions show varied results across models, with some struggling even with rich context, indicating that complex analytical tasks remain a challenge. ‘Funnel’ questions benefit significantly from context, especially for models like Gemini 1.5, suggesting that structured, narrowing inquiries are well-suited to context-enhanced responses. The ‘closed’ questions yield the highest accuracy for all models, regardless of context. In this regard, it is important to note that the revised evaluation metric for closed questions, which directly compares the model’s ‘yes’/‘no’ response to the ground truth in contrast to the RAGAS metric used for other question types.
5 Related Work
[…]
RAG for Long Documents RAG models offer a promising approach for enabling LLM s to search and extract relevant information from lengthy documents or extensive collections. The common strategy of splitting documents into smaller, more manageable chunks that fit within the LLM’s context window has its limitations, as highlighted by recent studies  Barnett et al. (2024). These limitations include the model’s failure to accurately extract answers even when they are present within the provided context, particularly when there is excessive noise or contradictory information. To address these challenges, researchers have proposed novel techniques. HippoRAG, a neurobiologically inspired long-term memory system designed for LLM s to handle long documents more effectively, aims to mitigate the limitations of current RAG models  Gutiérrez et al. (2024).  Gao et al. (2023) provide a comprehensive survey of RAG methods, categorizing them into three paradigms: Naive RAG, Advanced RAG, and Modular RAG. The authors highlight the remarkable adaptability of Modular RAG, which allows for module substitution or reconfiguration to address specific challenges, surpassing the fixed structures of Naive and Advanced RAG. Modular RAG integrates new modules or adjusts interaction flow among existing ones, enhancing its applicability across different tasks. The survey also discusses the concept of adaptive retrieval in RAG, exemplified by methods like Flare  Jiang et al. (2023b) and Self-RAG Asai et al. (2023). These approaches refine the RAG framework by enabling LLM s to actively determine the optimal moments and content for retrieval, enhancing the efficiency and relevance of the sourced information. Despite these advancements,  Gao et al. (2023) emphasize that further research is needed to fully understand the intricacies of applying RAG to long documents and to develop more robust and reliable methods.
6 Conclusion
@@
Our comprehensive analysis indicates that NEPA documents pose a significant challenge for LLM s, particularly in terms of understanding the complex semantics and effectively processing the lengthy documents. The findings reveal that models augmented with the RAG technique surpass those LLM s that are simply provided with the PDF content as long context. This suggests that incorporating more relevant knowledge retrieval processes can significantly enhance the performance of LLM s on complex document comprehension tasks like those found in the NEPA domain. In addition, we noticed that these LLM s struggle to use long input contexts to answer more difficult questions that require multiple steps of reasoning. For example, models performed best in answering closed questions and worst in answering divergent and problem-solving questions. However, when these models are augmented with additional context, as seen in the RAG setup, their performance can be significantly improved. We encourage the research community to focus on advancing the complementary use of long-context LLM s and RAG techniques together, rather than treating them as distinct, separate approaches.
7 Limitations
@@
Similar to other applications of LLMs, our proposed system for EIS long documents also has some limitations. We list those limitations as follows:
@@
Restriction of token limitation on full PDF context. While we are able to use the Gemini model with token length as 1.5 million, we could only use 128K tokens per query for response generation with Claude and GPT-4. Thus, we need to truncate the content of Full PDF to run these two LLM models. This might cause the performance drop on the context as Full PDF with questions from EIS documents. In future work, we should analyze more carefully about the impact of token truncation for Full PDF context.
@@
Uncertainty of generated responses by LLMs. Due to budget constraints, we conducted only one phase of response generation across different configurations. This introduces a risk of uncertain outputs, meaning that LLMs might generate different responses each time, even with the same input, as demonstrated in another study Wagle et al. (2024). In future work, we plan to run LLMs multiple times and analyze the effect of this uncertainty in response generation.
Challenges of human judgment. Currently, we leverage human evaluation as a preliminary proxy measure for qualitative analysis of benchmark. In future work, we plan to involve more NEPA experts in a more systematic manner to expand the dataset with human judgment results and to perform proper adjudication meetings between NEPA experts to reconcile conflicting results.
@@
Random errors when using RAGAS We noticed that RAGAS evaluation framework Es et al. (2023) produces random errors and "NaN" values in specific types of question. While running experiments to analyze the details of this fall outside of scope of this work, our preliminary tests lead us to assume that it might be due to the way the framework interacts with generative models and tries to conform their output to its formats.
@@
To address the potential bias in the answer correctness evaluation process, we assess both factual and semantic correctness in the evaluation. For semantic correctness, we utilize the BGE Xiao et al. (2023) as the embedding model and we calculate the semantic similarity between the model’s outputs and the ground-truth answers independently of GPT-4’s own evaluation mechanisms. By combining both factual and semantic correctness, we aim to accurately reflect the true performance of various models, including GPT-4.
8 Ethical Consideration
It has generally been the norm to assume that previously published work can be used as-is without having to consider the inherited ethical issues. However, in present times, researchers should not “simply assume that […] research will have a net positive impact on the world” Hecht et al. (2021). We acknowledge that this applies not just to new work, but also when using existing work in the way that we have done.
While working on this project, care has been taken to ensure that any and all data was anonymized and no Personally Identifiable Information (PII) is present in the data used. We had domain experts in the team throughout the process, thereby ensuring they were aware of all the potential risks and benefits.
While we do not anticipate the novel work presented here to introduce new ethical concerns in and by themselves, we do recognize that there may also be pre-existing concerns and issues of the data, models, and methodologies we have used for this paper. In particular, it has been seen that LLMs, like the ones used in this work, exhibit a wide variety of bias – religious, gender, race, profession, and cultural – and frequently generate answers that are incorrect, misogynistic, antisemitic, and generally toxic Abid et al. (2021); Buolamwini and Gebru (2018); Liang et al. (2021); Nadeem et al. (2021); Welbl et al. (2021). However, when used within the parameters of our experiments detailed in this paper, we did not see such behaviour from any of the models. To our knowledge, when used as intended, our models do not pose additional ethical concerns than any other LLM.
9 Acknowledgements
This work was supported by the Office of Policy, U.S. Department of Energy, and Pacific Northwest National Laboratory, which is operated by Battelle Memorial Institute for the U.S. Department of Energy under Contract DE-AC05–76RLO1830. This paper has been cleared by PNNL for public release as PNNL-SA-199339.
Appendix A Question Definitions

### v3

@@ 3.1. Supporting multiple LLM providers
No Context: Questions are presented to the models without any additional context. As such, one can think of this mode as analogous to using an LLM-based chatbot. While this strategy can work well with popular domains such as sport or literature, we assume that the NEPA domain may be challenging for the LLM s to provide accurate answers. In other words, for our benchmark, this setting can be considered a test of the LLM s’ ability to answer out-of-general-domain questions. As such, this setting is usually expected to return low performance.
@@
PDF Context: In addition to the question, we also provide the model the text from the entire PDF document from which the question was generated. Since we do not inform the model which part of the document to look at, the accuracy of generated responses will heavily depend on the models’ ability to pick the correct context from the very large scale textual information provided. We expect this setting to yield performance better than the no context scenario. This setup evaluates the models’ ability to process and extract relevant information from lengthy, complex documents.
@@
RAG Context:  Employing a RAG approach, this setup provides the top-$k$ most relevant text chunks retrieved from a vector database containing text chunks of NEPA documents. In our RAG model implementation, when a question is inputted for LLM generation, the corresponding context is extracted as a relevant passage from a given NEPA document. We use the standard RAG setup where we encode both question and retrieved passages with the BGE embedding model (Xiao et al., 2023). Cosine similarity scores are used to assess the similarity between the question and the contexts. In our case, the number of top-ranked relevant text chunks extracted is set at $k=3$. This scenario assesses the models’ performance when given highly relevant, focused information that is specifically retrieved to answer the question at hand.
@@
Gold Context:  In this configuration, we include the actual text excerpt from which the question was generated in the prompt, alongside the question content. This scenario represents an ideal case where the most relevant information is perfectly identified and provided. While the situation where users manually identify the correct text passage is rare in practice, we simulate this scenario to measure how well LLM s can perform if we were able to extract relevant passages with very high accuracy. This setup serves as an upper bound for performance, testing the models’ ability to comprehend and utilize perfectly relevant information.
3.3. Supporting custom prompt templates
We designed the prompt engineering system of MAPLE to be flexible and customizable through a template-based approach. We implemented a system where users can provide prompt templates through external files, with designated placeholders for question and context elements. The system automatically formats these templates during evaluation by replacing placeholders with actual content. We also included validation checks to ensure required placeholders are present for specific context types. For example, we require both question and context placeholders for gold context mode evaluations, while basic no-context evaluation requires only the placeholder for question. An example prompt template with placeholders for question and context is shown below. Through this design, we enabled users to experiment with different prompting strategies while ensuring consistent prompt structures across evaluations.
Prompt template with placeholder You are a helpful assistant who will answer questions about environmental policies from a context provided in the prompt.
Question: {placeholder for question}
Answer the question from the context.
Context: {placeholder for context}
3.4. Evaluation metrics
We leveraged the RAGAS evaluation framework (Es et al., 2023) to assess LLM responses against benchmark ground truth answers. We specifically focused on the answer correctness score, which is configured to combine two key components: factual correctness and semantic correctness. For factual correctness, which comprises $75\%$ of the final score, we utilized GPT-4 to evaluate the accuracy of generated answers at the phrase/clause level compared to ground truth answer in the benchmark. We complemented this with semantic correctness, weighted at $25\%$, where we employed the BGE embedding model (Xiao et al., 2023) to compare vector embeddings of predicted and expected answers. We chose this comprehensive evaluation approach over traditional overlap-based metrics like BLEU scores(Papineni et al., 2002), as it efficiently captures both factual accuracy and semantic similarity of the generated responses and ground truth answers.
@@ 3.4. Evaluation metrics
We employ a separate evaluation method for closed type questions which have only ‘yes’ or ‘no’ answers using the following steps. First, we process the responses generated by the LLM s to extract only the ‘yes’ or ‘no’ answer from the generated response. Any additional explanatory text or elaboration is removed from the response. Thereafter, the cleaned binary response is compared directly with the ground truth answer. If the cleaned response exactly matches the ground truth, it is assigned a full score of $100\%$. If there is any discrepancy, including cases where the model fails to provide a clear ”Yes” or ”No” answer, the response is scored as $0\%$. This straightforward evaluation method for closed questions allows us to assess the models’ ability to provide accurate binary responses without being influenced by any uncertainties introduced by the RAGAS evaluation framework. We refer this metric as answer correctness score for closed type questions.
4. Performance Analysis
In this section, we examine the performance of the LLM s on NEPAQuAD v1.0 using the MAPLE framework (as presented in Section 3). First, we compare the performance of the five frontier LLM sacross various contexts (Section 4.1). Then, we compare the model performance across various question types (Section 4.2).
4.1. Evaluating Different QA Contexts
@@
Table 5 reports the overall performance of the models across various QA contexts used in the evaluation. We observed that for the task with no context, the GPT-4 model produces the most accurate results by far. However, when PDF documents are provided as context, Claude surpassing GPT-4 and Gemini in the answer correctness. Despite the fact that Gemini is able to handle very long contexts (1.5M tokens), it is surprising to see its performance drop when provided with PDF documents as additional contexts. This may be due to the model struggling to reason over the large amount of relevant and irrelevant content in the EIS document. Open-source models (LLama3.1 and Mistral) generally underperform compared to closed LLM s across most context types, including PDF, RAG, and gold contexts. However, in the no-context scenario, all models exhibit similar performance, suggesting that the availability and quality of context significantly influence the performance gap between open-source and closed models.
@@
Table 5. Answer correctness of LLM responses on the NEPAQuAD v1.0 benchmark across different context types. The best-performing setting for each model is shown in bold.
@@
4.2. Evaluating Different Question Types
@@
Figure 4. The evaluation results meassured by the Answer Correctness scores of each LLM used with 4 scenarios of using context over each question types
@@
Figure 4 shows the performance of the models across different question types. The process, comparison and evaluation questions generally benefit from richer contexts, with closed-source models demonstrating a more significant improvement. Convergent and recall type questions prove challenging for all models, especially in no-context scenarios, highlighting the difficulty of synthesizing information or retrieving specific details without supporting context. Problem-solving questions show varied results across models, with some struggling even with rich context, indicating that complex analytical tasks remain a challenge. Funnel questions benefit significantly from context, especially for models like Gemini 1.5, suggesting that structured, narrowing inquiries are well-suited to context-enhanced responses. The closed questions yield the highest accuracy for all models, regardless of context. It is important to note that the evaluation metric for closed questions directly compares the model’s yes/no response to the ground truth, in contrast to the RAGAS metric used for other question types.
5. Related Works
Benchmarking LLMs for socio-scientific domains In recent years we have witnessed the emergence of numerous domain-specific benchmarks designed to rigorously evaluate LLM s across specialized fields. In the context of environment and environmental sciences, notable contributions include benchmark suite for water engineering (Yang et al., 2024), the ELLE dataset (Guo et al., 2025), the EnviroExam benchmark (Huang et al., 2024), and the WeQA benchmark (Meyur et al., 2024) focusing on RAG-based evaluation in the wind energy domain. Furthermore, specialized benchmarks for professional domains have also surfaced such as the LawBench benchmark (Chen et al., 2023), CFinBench (Guo et al., 2024), and TransportationGames (Zhang et al., 2024) to evaluate knowledge of LLM s in legal, financial and transport domains respectively. SciEval (Han et al., 2023) is another example which provides a multi-level evaluation benchmark for scientific research capabilities of LLMs. SciBench (Wang et al., 2023) and FrontierMath (Glazer et al., 2024) are premiere benchmarks to assess scientific and mathematical problem‐solving capabilities of LLM s. More recently, PhysReason (Zhang et al., 2025), QASA (Lee et al., 2023) and the CURIE dataset (Cui et al., 2025) have become popular benchmark dataset to evaluate LLM performance on scientific reasoning tasks.
Frameworks for LLM Evaluations The evaluation of LLM s has recently evolved beyond simple task-based metrics toward more sophisticated frameworks that address specific evaluation challenges across diverse domains. More recently, complementary approaches to tackle different aspects of LLM evaluation have emerged (Yao et al., 2024; Joty et al., 2024). Efforts to standardize evaluation processes have produced several notable toolkits, including Evalverse (Kim et al., 2024b), FMEval (Franceschi et al., 2024), and LLMeBench (Dalvi et al., 2023).For specialized RAG-based frameworks, Guinet et al. (2024) presents an automated evaluation protocol utilizing task‐specific exam generation. One of the most famous evaluation frameworks is LM Eval Harness (Gao et al., 2024), which provides a unified platform for measuring LLM performance across hundreds of standardized tasks and has become a cornerstone for reproducible model evaluation in the research community. Similarly, RAGAS (rag, 2023) provides multiple types of metrics to evaluate LLM responses on dimensions like faithfulness and context relevance, but does not provide a comprehensive workflow framework for end-to-end evaluation. DeepEval (dee, 2023) offers testing capabilities for LLM applications but is primarily optimized for OpenAI models, potentially limiting its applicability across different model providers. Our proposed MAPLE framework offers a transparent and modular approach that seamlessly integrates multiple LLM providers, supports customizable prompt templates, handles various context types, and incorporates diverse evaluation metrics within a unified evaluation pipeline.
6. Discussion and Conclusion
@@
Our analysis indicates that NEPA regulatory reasoning tasks pose a significant challenge for LLM s, particularly in terms of understanding the complex semantics and effectively processing the lengthy documents. The findings reveal that models augmented with the RAG technique surpass those that are simply provided with the PDF content as long context. This suggests that incorporating more relevant knowledge retrieval processes can significantly enhance the performance of LLM s on complex document comprehension tasks like those found in the NEPA domain. In addition, we noticed that these models struggle to use long input contexts to answer more difficult questions that require multiple steps of reasoning. For example, models performed best in answering closed questions and worst in answering divergent and problem-solving questions. However, when these models are augmented with additional context, as seen in the RAG setup, their performance can be significantly improved.
7. Limitations
@@
As with all work, our proposed system for EIS long documents also has some limitations. We list those limitations as follows:
@@
Restriction of token limitation on full PDF context. While we are able to use the Gemini model with token length as 1.5 million, we could only use 128K tokens per query for response generation with Claude and GPT-4. Thus, we needed to truncate the content of Full PDF to run these two models. This might cause the performance drop on the full PDF context mode. In future work, we should analyze more carefully about the impact of token truncation.
@@
Uncertainty of generated responses by LLM s. Due to budget constraints, we conducted only one phase of response generation across different configurations. This introduces a risk of uncertain outputs, as LLMs might generate different responses each time, even with the same input, as demonstrated in another study (Wagle et al., 2024). Multiple runs of the same model might mitigate this effect.
Challenges of human judgment. Currently, we leverage human evaluation only on a subset of entries in the benchmark. With sufficient time and resources, a thorough evaluation of the entire dataset following the same methodology will likely make the dataset even more robust.
Imbalance of question types. In the current benchmark, just around half the questions are closed type. This is due to a myriad of reasons, including difficulty in generating open questions and the time-consuming nature of open question evaluation. While open-questions do take significantly more human review and effort, a future benchmark with even more such questions will undoubtedly provide an even tougher challenge to LLMs.
@@
Random errors when using RAGAS We noticed that RAGAS evaluation framework (Es et al., 2023) produces random errors and ”NaN” values in specific types of question. While running experiments to analyze the details of this fall outside of scope of this work, our preliminary tests lead us to assume that it might be due to the way the framework interacts with generative models and tries to conform their output to its formats.
@@
To address the potential bias in the answer correctness evaluation process, we assess both factual and semantic correctness in the evaluation. For semantic correctness, we utilize the BGE (Xiao et al., 2023) as the embedding model and we calculate the semantic similarity between the model’s outputs and the ground-truth answers independently of GPT-4’s own evaluation mechanisms. By combining both factual and semantic correctness, we aim to accurately reflect the true performance of various models, including GPT-4.
Acknowledgement
This work was supported by the Office of Policy, U.S. Department of Energy, and Pacific Northwest National Laboratory, which is operated by Battelle Memorial Institute for the U.S. Department of Energy under Contract DE-AC05–76RLO1830. This paper has been cleared by PNNL for public release as PNNL-SA-211465.
Appendix A Prompt Template
The template for the prompt is shown below:
Prompt template with placeholder 1. Generate {placeholder for number of questions} {placeholder for type of question} questions from the following context.
Context: {placeholder for gold passage}
2. Definition of {placeholder for type of question} question: {placeholder for definition provided by NEPA experts}
3. The questions can be similar to following sample questions:
placeholder for sample questions provided by NEPA experts
4. Return the questions in CSV format with columns question / answer / proof.
We provided appropriate placeholders for the type and number of questions to be generated, the excerpt from which the question will be generated along with the sample questions of the same type as provided by the NEPA experts. We restricted the output for each prompt in a CSV format with three fields: question, answer, and proof. The ‘proof’ column stored the part of the gold passage that the model picked as evidence for the provided answer to the question.
Appendix B Question Definitions